OpenStudy (anonymous):

BREAKING NEWS!: i need someone to check my answers on 3 problems but no one seems to know how to help

myininaya (myininaya):

It is hard to read if you don't separate your numerator from your denominator properly. I think I can translate though. I think you mean (4x^2 + x - 7) / (x^2 - 7) + (4x - 3x^2) / (x^2 - 7)

myininaya (myininaya):

Since the denominators are the same you can now add the numerators (and yes I said add because you have a plus sign in between the fractions)

myininaya (myininaya):

\[(4x^2+x-7)+(4x-3x^2) =? \]

myininaya (myininaya):

gather your like terms

OpenStudy (anonymous):

yeah sorry I didn't put them in parenthesis

myininaya (myininaya):

\[4x^2-3x^2=? \] \[x+4x=? \]

OpenStudy (anonymous):

So number 1 is B then...

myininaya (myininaya):

no

myininaya (myininaya):

4-3=1 and 1+4=5